translate to a algebratic equation then solve.
the width of a rectangular playground is 18 meters less thsn the lenght. find the width when the perimeter of the rectangle is 84 meters.

Answer by rfer(16322)

p=2L+2W
L=x
W=x-18
2x+2(x-18)=84
2x+2x-36=84
4x=120
x=30
proof
60+24=84